# The Angular Momentum Penrose Inequality

### v3

@@ 4 Stage 1: Axisymmetric Jang Equation
4.1 Function Spaces and Regularity Framework
We first establish the precise function spaces required for rigorous analysis.
@@ 4.1 Function Spaces and Regularity Framework
For $k\in\mathbb{N}_{0}$, $\beta\in(0,1)$, and weight $\tau\in\mathbb{R}$, define the weighted Hölder space on an asymptotically flat manifold $(M,g)$ with asymptotic radial coordinate $r(x):=|x|$ in the end:
@@
where the norm is:
$\|u\|_{C^{k,\beta}_{-\tau}}:=\sum_{|\beta|\leq k}\sup_{x\in M}\langle r(x)\rangle^{\tau+|\beta|}|D^{\beta}u(x)|+[D^{k}u]_{\beta,-\tau-k-\beta},$
with $\langle r\rangle:=(1+r^{2})^{1/2}$ (the Japanese bracket), and the weighted Hölder seminorm:
$v]_{\beta,\delta}:=\sup_{\begin{subarray}{c}x\neq y\\ d(x,y)<\mathrm{inj}(M)/2\end{subarray}}\min(\langle r(x)\rangle,\langle r(y)\rangle)^{-\delta}\frac{|v(x)-v(y)|}{d(x,y)^{\beta}}.$
Here $\mathrm{inj}(M)$ denotes the injectivity radius. A function $u\in C^{k,\beta}_{-\tau}(M)$ satisfies $|u(x)|=O(r^{-\tau})$ as $r\to\infty$.
This follows the conventions of Bartnik [9] and Lockhart–McOwen [35]. The choice $\tau>1/2$ in Definition 4.2 ensures finite ADM mass.
@@
Initial data $(M,g,K)$ is asymptotically flat with decay rate $\tau>1/2$ if there exists a compact set $K_{0}\subset M$ and a diffeomorphism $\Phi:M\setminus K_{0}\to\mathbb{R}^{3}\setminus\overline{B_{R}}$ for some $R>0$, such that in the coordinates $x=\Phi(p)$:
  1. (AF1)
    Metric decay: $g_{ij}-\delta_{ij}\in C^{2,\beta}_{-\tau}(M\setminus K_{0})$, i.e.,
    $|g_{ij}(x)-\delta_{ij}|\leq C|x|^{-\tau},\quad|\partial_{k}g_{ij}(x)|\leq C|x|^{-\tau-1},\quad|\partial_{k}\partial_{\ell}g_{ij}(x)|\leq C|x|^{-\tau-2};$
  2. (AF2)
    Extrinsic curvature decay: $K_{ij}\in C^{1,\beta}_{-\tau-1}(M\setminus K_{0})$, i.e.,
    $|K_{ij}(x)|\leq C|x|^{-\tau-1},\quad|\partial_{k}K_{ij}(x)|\leq C|x|^{-\tau-2};$
  3. (AF3)
    Finite ADM mass: The ADM mass, defined by the limit
    $M_{\mathrm{ADM}}:=\lim_{R\to\infty}\frac{1}{16\pi}\oint_{S_{R}}(\partial_{j}g_{ij}-\partial_{i}g_{jj})\nu^{i}\,dA,$
    exists and is finite. Here $S_{R}=\{|x|=R\}$ and $\nu=x/|x|$ is the Euclidean outward normal.
The condition $\tau>1/2$ ensures convergence of the ADM integral: the integrand is $O(R^{-\tau-1})$, so the surface integral is $O(R^{2-\tau-1})=O(R^{1-\tau})\to 0$ as $R\to\infty$ when $\tau>1$; the weaker condition $\tau>1/2$ suffices by more refined analysis using the constraint equations (see [9, Theorem 4.2]).
@@
Initial data $(M,g,K)$ satisfies the dominant energy condition (DEC) if:
$\mu\geq|\boldsymbol{j}|_{g},\quad\text{where }\mu=\frac{1}{2}(R_{g}+(\mathrm{tr}_{g}K)^{2}-|K|_{g}^{2}),\quad\boldsymbol{j}_{i}=D^{k}K_{ki}-D_{i}(\mathrm{tr}_{g}K).$
Here $\mu$ is the energy density and $\boldsymbol{j}$ is the momentum density vector field (see Remark 1.9). For vacuum data ($\mu=|\boldsymbol{j}|_{g}=0$), DEC is automatic.
@@
A closed surface $\Sigma\subset M$ is a marginally outer trapped surface (MOTS) if the outward null expansion vanishes: $\theta^{+}:=H_{\Sigma}+\mathrm{tr}_{\Sigma}K=0$, where $H_{\Sigma}=\mathrm{div}_{\Sigma}(\nu)$ is the mean curvature (trace of the second fundamental form with respect to the outward normal $\nu$), and $\mathrm{tr}_{\Sigma}K:=K_{ij}(\delta^{ij}-\nu^{i}\nu^{j})$ is the trace of $K$ restricted to $\Sigma$. The surface is outermost if no other MOTS encloses it, i.e., lies in the exterior region $M\setminus\overline{\mathrm{Int}(\Sigma)}$.
A MOTS is stable if the principal eigenvalue of the MOTS stability operator
$L_{\Sigma}:W^{2,2}(\Sigma)\to L^{2}(\Sigma),\quad L_{\Sigma}[\psi]:=-\Delta_{\Sigma}\psi-\left(|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu)\right)\psi-\mathrm{div}_{\Sigma}(X\psi)-X\cdot\nabla_{\Sigma}\psi$ (14)
satisfies $\lambda_{1}(L_{\Sigma})\geq 0$. Here:
  • $A_{\Sigma}$ is the second fundamental form of $\Sigma$ in $(M,g)$, with $|A_{\Sigma}|^{2}=\sum_{i,j}(A_{ij})^{2}$;
  • $\mathrm{Ric}_{g}(\nu,\nu)=R_{ij}\nu^{i}\nu^{j}$ is the Ricci curvature in the normal direction;
  • $X:=(K(\nu,\cdot))^{\top}\in\Gamma(T\Sigma)$ is the tangential projection of $K(\nu,\cdot)$ to $\Sigma$, i.e., $X^{i}=K_{j}{}^{i}\nu^{j}-K_{jk}\nu^{j}\nu^{k}\nu^{i}$.
Since the first-order terms make $L_{\Sigma}$ non-self-adjoint, the principal eigenvalue $\lambda_{1}(L_{\Sigma})$ is defined as:
$\lambda_{1}(L_{\Sigma}):=\inf\{\Re(\lambda):\lambda\in\sigma(L_{\Sigma})\},$
where $\sigma(L_{\Sigma})\subset\mathbb{C}$ is the spectrum. By the Krein–Rutman theorem [33] applied to the formal adjoint, there exists a real eigenvalue achieving this infimum with a positive eigenfunction.
For time-symmetric data ($K=0$), we have $X=0$ and the operator simplifies to the self-adjoint form $L_{\Sigma}[\psi]=-\Delta_{\Sigma}\psi-(|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu))\psi$, for which the variational characterization $\lambda_{1}=\inf_{\|\psi\|_{L^{2}}=1}\langle L_{\Sigma}\psi,\psi\rangle_{L^{2}}$ applies.
This definition follows Andersson–Mars–Simon [6] and Andersson–Metzger [7].
Remark 4.5 (Strictly Stable MOTS and Cylindrical Decay Rate).
The hypothesis of strict stability ($\lambda_{1}(L_{\Sigma})>0$) in Theorem 1.2 is directly connected to the cylindrical end decay rate $\beta_{0}$ in the Jang construction (Theorem 4.12):
  1. (i)
    Spectral correspondence: By [7, Proposition 3.4], the cylindrical decay rate satisfies $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}$ for strictly stable MOTS. This relationship arises from the linearized Jang equation at the MOTS.
  2. (ii)
    Decay rate implications: For $\lambda_{1}(L_{\Sigma})>0$:
    • The Jang metric converges exponentially to the cylinder: $\bar{g}=dt^{2}+g_{\Sigma}+O(e^{-\beta_{0}t})$;
    • The decay rate $\beta_{0}>0$ ensures Fredholm theory applies with weight $\beta\in(-\beta_{0}/2,0)$;
    • All geometric quantities ($R_{\bar{g}}$, $\Lambda_{J}$, etc.) decay exponentially along the cylindrical end.
  3. (iii)
    Marginally stable case: For $\lambda_{1}(L_{\Sigma})=0$, a limiting argument using subleading spectral terms gives $\beta_{0}=2$ (see Lemma 4.18, Step 4). The proof extends to this case with minor modifications to the weighted space analysis.
  4. (iv)
    Physical interpretation: Strictly stable MOTS represent “isolated” horizons that are dynamically stable under small perturbations. The spectral gap $\lambda_{1}>0$ quantifies the “stiffness” of the horizon against deformations. Marginally stable MOTS (e.g., at the threshold of black hole formation) have $\lambda_{1}=0$.
The hypothesis (H4) in Theorem 1.2 requires $\lambda_{1}(L_{\Sigma})>0$, which is satisfied by generic black hole data and, in particular, by all sub-extremal Kerr slices.
@@ A MOTS is stable if the principal eigenvalue of the MOTS stability operator
Let $(M,g,K)$ be asymptotically flat, axisymmetric initial data satisfying DEC with Killing field $\eta=\partial_{\phi}$ and axis $\Gamma=\{\eta=0\}$. Let $\Sigma$ be a strictly stable outermost MOTS. Then:
  1. (i)
    $\Sigma$ has spherical topology: $\Sigma\cong S^{2}$ (by the Galloway–Schoen theorem [27]).
  2. (ii)
    $\Sigma$ intersects the axis $\Gamma$ at exactly two points (the “poles”): $\Sigma\cap\Gamma=\{p_{N},p_{S}\}$.
  3. (iii)
    Away from the poles, the orbit radius is strictly positive: $\rho|_{\Sigma\setminus\{p_{N},p_{S}\}}>0$.
  4. (iv)
    The orbit radius vanishes linearly at the poles: $\rho(x)=O(\mathrm{dist}_{g}(x,p_{\pm}))$ as $x\to p_{\pm}$.
@@
Let $(M,g,K)$ be asymptotically flat, axisymmetric initial data satisfying DEC, and let $\Sigma$ be a stable outermost MOTS with poles $p_{N},p_{S}=\Sigma\cap\Gamma$. Let $\mathcal{T}[\bar{f}]$ be the twist perturbation term (24) in the orbit-space Jang equation. Then:
  1. (i)
    Twist scaling at poles: Near each pole $p\in\{p_{N},p_{S}\}$:
    $|\mathcal{T}[\bar{f}](x)|\leq C\cdot\rho(x)^{2}\cdot|\bar{\nabla}\bar{f}|(x)\leq C^{\prime}\cdot d(x,p)^{2}\quad\text{as }x\to p,$ (15)
    where $d(x,p)=\mathrm{dist}_{g}(x,p)$ is the distance to the pole.
  2. (ii)
    Integrability: The twist term is integrable over $\Sigma$ with respect to the induced area measure:
    $\int_{\Sigma}|\mathcal{T}[\bar{f}]|\,dA_{\Sigma}<\infty.$ (16)
  3. (iii)
    Perturbative control: The twist contribution to the Jang operator remains uniformly bounded:
    $\sup_{x\in\Sigma}|\mathcal{T}[\bar{f}](x)|\leq C_{\mathcal{T}}<\infty,$ (17)
    where $C_{\mathcal{T}}$ depends only on the initial data.
In particular, the presence of poles where $\rho=0$ does not obstruct the Jang existence theory.
@@
$|\mathcal{T}[\bar{f}](x)|\leq C\cdot\rho(x)^{2}\cdot|\omega(x)|\cdot|\bar{\nabla}\bar{f}|(x)=O(d(x,p)^{2}\cdot 1\cdot O(1))=O(d(x,p)^{2}).$
@@
The coordinate singularity at the rotation axis $\Gamma=\{r=0\}$ in Weyl–Papapetrou coordinates requires careful treatment in the weighted Hölder space framework. Specifically:
  1. (i)
    Coordinate singularity vs. geometric regularity: Although the metric coefficient $g_{\phi\phi}=\rho^{2}\to 0$ as $r\to 0$, this reflects the coordinate choice rather than a geometric singularity. The manifold $(M,g)$ is smooth across the axis, and axis regularity conditions (AR1)–(AR3) ensure that tensor fields (including the twist potential $\omega$) extend smoothly when expressed in Cartesian-like coordinates near the axis.
  2. (ii)
    Weighted norms and the axis: The weighted Hölder norm $\|\cdot\|_{C^{k,\beta}_{-\tau}}$ (Definition 4.1) involves the radial weight $\langle r\rangle^{-\tau}$ for asymptotic decay, but near the axis we use the $\rho$-weighted regularity $C^{k,\beta}_{\rho}$ as in condition (AR3). This hybrid weighting—polynomial in $r$ for asymptotics, $\rho$-scaled for the axis—is standard in the analysis of axisymmetric elliptic problems [19, 23].
  3. (iii)
    Elliptic regularity at the axis: The Jang operator and AM-Lichnerowicz operator, when reduced to the orbit space $\mathcal{Q}$, become degenerate elliptic at the axis (the coefficient of $\partial_{r}^{2}$ vanishes like $r^{2}$ in certain formulations). Standard regularity theory [40] for such edge-degenerate operators ensures that solutions inherit the axis regularity of the data, provided conditions (AR1)–(AR3) hold. The key point is that the twist potential $\omega$ satisfying (AR1)–(AR2) produces twist perturbation terms $\mathcal{T}$ that remain in the appropriate weighted space.
@@
Step 1: Corner geometry. Near a corner point $\bar{p}\in\bar{\Sigma}\cap\partial\mathcal{Q}$, introduce local coordinates $(\sigma,\tau)$ where:
  • $\sigma\geq 0$ is the distance to the axis $\partial\mathcal{Q}$;
  • $\tau$ is arc-length along $\partial\mathcal{Q}$, with $\tau=0$ at the corner.
The MOTS curve $\bar{\Sigma}$ meets $\partial\mathcal{Q}$ at an angle $\theta_{0}\in(0,\pi)$, so near the corner:
@@ 4.3 Axisymmetric Setting
Theorem 4.12 relies on the twist decay estimate established in Step 2c of the proof: the twist perturbation term $\mathcal{T}[f]$ satisfies $|\mathcal{T}[f]|=O(s)$ as $s\to 0$, where $s=\mathrm{dist}(\cdot,\Sigma)$.
The principal Jang operator terms scale as $O(s^{-1})$ near the MOTS. For the perturbation argument (Lemma 4.16) to succeed, the twist must be subdominant. The estimate $|\mathcal{T}|=O(s)$ provides a factor of $s^{2}$ separation.
Physical intuition for twist boundedness: The twist 1-form $\omega$ encodes frame-dragging effects. For stable MOTS in axisymmetric geometries, the twist potential satisfies an elliptic equation ($d\omega=0$ combined with regularity conditions at the axis) that enforces boundedness. The key physical insight is that:
  • Frame-dragging is maximal on the horizon, not as one approaches it from outside;
  • The axis regularity conditions (requiring $\omega\to 0$ smoothly as $\rho\to 0$) prevent twist from diverging at the poles;
  • The elliptic equation propagates this boundary regularity throughout the domain.
The mathematical manifestation is that $|\omega|\leq C_{\omega,\infty}<\infty$ uniformly on $\mathcal{Q}$, including near the MOTS.
Verification checkpoints:
  1. (V1)
    Twist 1-form regularity: The axis regularity conditions (AR1)–(AR3) in Remark 4.10 imply $|\omega|\leq C_{\omega,\infty}<\infty$ on the orbit space $\mathcal{Q}$. This follows from standard elliptic theory for the twist potential equation.
  2. (V2)
    Orbit radius scaling: The twist term (24) contains a factor of $\rho^{2}$ in the numerator. Since $\rho$ is bounded on the compact MOTS $\Sigma$, this contributes a bounded factor.
  3. (V3)
    Gradient scaling: The denominator $\sqrt{1+|\nabla f|^{2}}=O(s^{-1})$ due to the logarithmic blow-up $f=C_{0}\ln s^{-1}+O(1)$. This factor produces the $O(s)$ decay.
  4. (V4)
    Pole regularity: At the poles $p_{N},p_{S}$ where $\rho=0$, the twist term vanishes: $\mathcal{T}(p_{\pm})=0$ (Lemma 4.8).
  5. (V5)
    Frame-dragging cancellation structure: The specific contraction structure in $\mathcal{T}[f]$ (involving the graph normal $\nu$) does not amplify frame-dragging effects near the horizon. The orbit-space normal $\bar{\nu}$ remains bounded as $s\to 0$.
Summary of Twist Scaling Mechanisms: The twist decay estimate $|\mathcal{T}[f]|=O(s)$ relies on three independent mechanisms:
  • Twist potential boundedness: $|\omega|\leq C_{\omega,\infty}$ (from elliptic PDE on orbit space);
  • Orbit radius prefactor: $\rho^{2}=O(1)$ (from dimensional reduction);
  • Gradient normalization: $\sqrt{1+|\nabla f|^{2}}=O(s^{-1})$ (from Jang blow-up theory).
Combined, these yield $|\mathcal{T}|=O(s)$.
@@
$\mathcal{J}_{\text{axi}}[\bar{f}]:=\bar{g}^{ij}\left(\frac{\bar{\nabla}_{ij}\bar{f}}{\sqrt{1+|\bar{\nabla}\bar{f}|^{2}}}-\bar{K}_{ij}\right)-\frac{\bar{f}^{i}\bar{f}^{j}}{1+|\bar{\nabla}\bar{f}|^{2}}\left(\frac{\bar{\nabla}_{ij}\bar{f}}{\sqrt{1+|\bar{\nabla}\bar{f}|^{2}}}-\bar{K}_{ij}\right)-\mathcal{T}[\bar{f}],$
@@
Let $(M^{3},g,K)$ be asymptotically flat, axisymmetric initial data with a stable outermost MOTS $\Sigma$. Let $s=\mathrm{dist}(\cdot,\Sigma)$ denote the signed distance to $\Sigma$, and let $\mathcal{T}[f]$ be the twist perturbation term (24) in the axisymmetric Jang equation. Then there exist constants $C_{\mathcal{T}}>0$ and $s_{0}>0$ depending only on the initial data such that:
@@
More precisely, $C_{\mathcal{T}}=C_{\omega,\infty}\cdot\rho_{\max}^{2}/C_{0}$, where:
  • $C_{\omega,\infty}=\sup_{\mathcal{Q}}|\omega|$ is the $L^{\infty}$ bound on the twist 1-form;
  • $\rho_{\max}=\sup_{x\in\Sigma}\rho(x)$ is the maximum orbit radius on $\Sigma$;
  • $C_{0}>0$ is the leading coefficient in the Jang blow-up $f=C_{0}\ln s^{-1}+O(1)$.
Scaling comparison: Since the principal Jang terms scale as $O(s^{-1})$ near $\Sigma$, while the twist term scales as $O(s)$, the twist is subdominant by a factor of $s^{2}$. This ensures that twist does not disrupt the blow-up analysis, preserving the cylindrical end structure required for the proof.
Critical observation: The constant $C_{\mathcal{T}}$ depends only on the initial data $(g,K)$ and the blow-up coefficient $C_{0}=|\theta^{-}|/2$, which is determined by the MOTS geometry. In particular:
  1. (a)
    $C_{\mathcal{T}}$ does not depend on higher derivatives $\nabla^{k}f$ for $k\geq 2$, which blow up as $O(s^{-k})$;
  2. (b)
    The twist term $\mathcal{T}[f]$ involves no second derivatives of $f$, only $f$ and $\nabla f$;
  3. (c)
    The bound holds uniformly for any function with logarithmic blow-up $f=C_{0}\ln s^{-1}+O(1)$;
  4. (d)
    At the poles $p_{N},p_{S}$ where $\Sigma$ intersects the axis, $\mathcal{T}(p_{\pm})=0$ since $\rho(p_{\pm})=0$ (Lemma 4.8).
See Appendix D for the complete verification that the twist does not alter the existence or character of the Jang solution.
Non-circularity verification: The argument above is not circular. To see this explicitly:
  1. (NC1)
[…]
  2. (NC2)
    The twist bound $|\omega|\leq C_{\omega,\infty}$ follows from elliptic regularity applied to the twist potential equation on the orbit space, which is determined by the initial data $(g,K)$ alone.
  3. (NC3)
    The orbit radius $\rho_{\max}=\sup_{\Sigma}\rho$ is a geometric quantity of the MOTS in the initial data.
  4. (NC4)
    The scaling $|\mathcal{T}[f]|=O(s)$ uses only that $f$ has logarithmic blow-up with some coefficient $C_{0}>0$, not the specific value. Thus, the estimate holds for any candidate solution in the iteration scheme of Lemma 4.16.
The logical flow is: initial data $\to$ MOTS geometry $\to$ blow-up coefficient $C_{0}$ and twist bound $C_{\omega,\infty}$ $\to$ perturbation estimate $|\mathcal{T}|\leq C_{\mathcal{T}}s$ $\to$ Jang existence via fixed-point argument. At no point does the constant $C_{\mathcal{T}}$ depend on the solution $f$ being constructed.
Remark 4.15 (Verification of Non-Circularity).
We verify that the fixed-point argument in Lemma 4.16 is non-circular by explicitly tracking the data dependencies of all constants.
Given data (independent of Jang solution):
  • $(M,g,K)$: initial data set with metric $g$ and extrinsic curvature $K$;
  • $\Sigma\subset M$: stable MOTS with mean curvature $H_{\Sigma}$ and $\mathrm{tr}_{\Sigma}K$;
  • $\eta$: axial Killing field with $\mathcal{L}_{\eta}g=\mathcal{L}_{\eta}K=0$.
A priori constants (computable from given data before solving Jang):
  1. (C1)
    $C_{0}=|\theta^{-}|/2=|H_{\Sigma}-\mathrm{tr}_{\Sigma}K|/2$: determined by the expansion of null geodesics orthogonal to $\Sigma$, computed entirely from $(g|_{\Sigma},K|_{\Sigma})$.
  2. (C2)
    $\rho_{\max}=\sup_{p\in\Sigma}|\eta|_{g}(p)$: the maximum orbit radius on the MOTS, a geometric quantity of $(M,g,\eta,\Sigma)$.
  3. (C3)
    $C_{\omega,\infty}$: the $L^{\infty}$ bound on the twist 1-form $\omega$, obtained by solving the linear elliptic equation $d\omega=0$, $\delta\omega=\star(d\eta^{\flat}\wedge\eta^{\flat})/\rho^{2}$ on the orbit space $\mathcal{Q}=M/\mathrm{U}(1)$. This depends only on $(g,K,\eta)$.
  4. (C4)
    $\lambda_{0}(\Sigma)$: the principal eigenvalue of the MOTS stability operator on $\Sigma$, determined by $(g|_{\Sigma},K|_{\Sigma})$.
Fixed-point iteration: The contraction map $\Phi:B_{\delta}\to B_{\delta}$ in Step 4 of Lemma 4.16’s proof is defined by:
$\Phi(v)=-L_{0}^{-1}\bigl(N[v]+\mathcal{T}[f_{0}+v]\bigr),$
where $f_{0}$ is the reference solution to the twist-free Jang equation, $L_{0}$ is the linearization, and $N[v]$ is the nonlinear remainder.
Key observation: The twist bound $|\mathcal{T}[f]|\leq C_{\mathcal{T}}s$ in hypothesis (P3) requires only $|\omega|\leq C_{\omega,\infty}$, $\rho\leq\rho_{\max}$, and $|\nabla f|\geq C_{0}/s-O(1)$. The last condition holds uniformly for all elements of $B_{\delta}$.
Conclusion: The constant $C_{\mathcal{T}}=C_{\omega,\infty}\cdot\rho_{\max}^{2}/C_{0}$ is computed from (C1)–(C3), which are determined before the Jang equation is solved. The fixed-point argument then constructs the solution $f=f_{0}+v$ without any circular dependence.

### v4

@@ 4 Stage 1: Axisymmetric Jang Equation
4.1 Function Spaces
We work with weighted Hölder spaces on asymptotically flat manifolds, following Bartnik [9] and Lockhart–McOwen [35].
@@ 4.1 Function Spaces
For $k\in\mathbb{N}_{0}$, $\beta\in(0,1)$, and weight $\tau\in\mathbb{R}$, define
@@
where $\|u\|_{C^{k,\beta}_{-\tau}}$ incorporates factors of $\langle r\rangle^{\tau+|\beta|}$ for derivatives $D^{\beta}u$, with $\langle r\rangle=(1+r^{2})^{1/2}$. Functions in $C^{k,\beta}_{-\tau}(M)$ decay like $O(r^{-\tau})$ at infinity.
@@
Initial data $(M,g,K)$ is asymptotically flat with decay rate $\tau>1/2$ if, outside a compact set, coordinates exist in which $g_{ij}-\delta_{ij}\in C^{2,\beta}_{-\tau}$ and $K_{ij}\in C^{1,\beta}_{-\tau-1}$, and the ADM mass
$M_{\mathrm{ADM}}:=\lim_{R\to\infty}\frac{1}{16\pi}\oint_{S_{R}}(\partial_{j}g_{ij}-\partial_{i}g_{jj})\nu^{i}\,dA$
is finite. The condition $\tau>1/2$ ensures convergence of this integral [9].
@@
The dominant energy condition (DEC) is $\mu\geq|\boldsymbol{j}|_{g}$, where the energy density is $\mu=\frac{1}{2}(R_{g}+(\mathrm{tr}_{g}K)^{2}-|K|_{g}^{2})$ and $\boldsymbol{j}_{i}=D^{k}K_{ki}-D_{i}(\mathrm{tr}_{g}K)$ is the momentum density.
@@
A closed surface $\Sigma$ is a marginally outer trapped surface (MOTS) if $\theta^{+}:=H_{\Sigma}+\mathrm{tr}_{\Sigma}K=0$. It is outermost if no MOTS encloses it, and stable if the principal eigenvalue of the stability operator
$L_{\Sigma}[\psi]:=-\Delta_{\Sigma}\psi-(|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu))\psi-\mathrm{div}_{\Sigma}(X\psi)-X\cdot\nabla_{\Sigma}\psi$
satisfies $\lambda_{1}(L_{\Sigma})\geq 0$, where $X=(K(\nu,\cdot))^{\top}$ is the tangential projection of $K(\nu,\cdot)$. We call the MOTS strictly stable if $\lambda_{1}(L_{\Sigma})>0$. This definition follows Andersson–Mars–Simon [6].
Remark 4.5.
Strict stability ($\lambda_{1}(L_{\Sigma})>0$) determines the exponential decay rate $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}$ in the Jang cylindrical end [7]. The spectral gap ensures Fredholm theory applies with appropriate weights.
@@
Let $(M,g,K)$ be asymptotically flat, axisymmetric initial data satisfying DEC with Killing field $\eta=\partial_{\phi}$ and axis $\Gamma=\{\eta=0\}$. Let $\Sigma$ be a strictly stable outermost MOTS. Then $\Sigma$ has spherical topology ($\Sigma\cong S^{2}$, by the Galloway–Schoen theorem [27]), $\Sigma$ intersects the axis $\Gamma$ at exactly two points (the “poles”): $\Sigma\cap\Gamma=\{p_{N},p_{S}\}$, away from the poles the orbit radius is strictly positive: $\rho|_{\Sigma\setminus\{p_{N},p_{S}\}}>0$, and the orbit radius vanishes linearly at the poles: $\rho(x)=O(\mathrm{dist}_{g}(x,p_{\pm}))$ as $x\to p_{\pm}$.
@@
Remark 4.8 (Twist Perturbation and Axis Regularity).
The estimate $|\mathcal{T}[\bar{f}]|=O(s)$ (where $s$ is distance to the MOTS) is crucial for the blow-up analysis. Near the rotation axis $\Gamma$, where the orbit radius $\rho\to 0$, one might worry that the twist term (involving $\rho^{-4}$) could be singular. However, as shown in Lemma 4.9, the twist term vanishes at the poles because the angular momentum potential $\omega$ and its derivatives scale with sufficiently high powers of $\rho$ (specifically $\partial\omega\sim\rho^{3}$ near the axis). Our function space framework (Appendix F) explicitly handles the weight degeneration at the axis, ensuring that the twist perturbation remains a lower-order term in the weighted Hölder norms, even at the poles.
@@
Let $(M,g,K)$ be asymptotically flat, axisymmetric initial data satisfying DEC, and let $\Sigma$ be a stable outermost MOTS with poles $p_{N},p_{S}=\Sigma\cap\Gamma$. Let $\mathcal{T}[\bar{f}]$ be the twist perturbation term (11) in the orbit-space Jang equation. Then: (i) the twist scales as $|\mathcal{T}[\bar{f}](x)|\leq C\cdot\rho(x)^{2}\cdot|\bar{\nabla}\bar{f}|(x)\leq C^{\prime}\cdot d(x,p)^{2}$ near each pole $p\in\{p_{N},p_{S}\}$, where $d(x,p)=\mathrm{dist}_{g}(x,p)$; (ii) the twist term is integrable over $\Sigma$ with respect to the induced area measure, $\int_{\Sigma}|\mathcal{T}[\bar{f}]|\,dA_{\Sigma}<\infty$; and (iii) the twist contribution to the Jang operator remains uniformly bounded, $\sup_{x\in\Sigma}|\mathcal{T}[\bar{f}](x)|\leq C_{\mathcal{T}}<\infty$, where $C_{\mathcal{T}}$ depends only on the initial data. In particular, the presence of poles where $\rho=0$ does not obstruct the Jang existence theory.
@@
$|\mathcal{T}[\bar{f}](x)|\leq C\cdot\rho(x)^{2}\cdot|\omega(x)|\cdot|\bar{\nabla}\bar{f}|(x)=O(d(x,p)^{2}\cdot 1\cdot O(1))=O(d(x,p)^{2}).$ (5)
@@
The coordinate singularity at the rotation axis $\Gamma=\{r=0\}$ in Weyl–Papapetrou coordinates requires careful treatment in the weighted Hölder space framework. Although the metric coefficient $g_{\phi\phi}=\rho^{2}\to 0$ as $r\to 0$, this reflects the coordinate choice rather than a geometric singularity; the manifold $(M,g)$ is smooth across the axis, and axis regularity conditions (AR1)–(AR3) ensure that tensor fields (including the twist potential $\omega$) extend smoothly when expressed in Cartesian-like coordinates near the axis. The weighted Hölder norm $\|\cdot\|_{C^{k,\beta}_{-\tau}}$ (Definition 4.1) involves the radial weight $\langle r\rangle^{-\tau}$ for asymptotic decay, but near the axis we use the $\rho$-weighted regularity $C^{k,\beta}_{\rho}$ as in condition (AR3). This hybrid weighting—polynomial in $r$ for asymptotics, $\rho$-scaled for the axis—is standard in the analysis of axisymmetric elliptic problems [19, 23]. The Jang operator and AM-Lichnerowicz operator, when reduced to the orbit space $\mathcal{Q}$, become degenerate elliptic at the axis (the coefficient of $\partial_{r}^{2}$ vanishes like $r^{2}$ in certain formulations). Standard regularity theory [40] for such edge-degenerate operators ensures that solutions inherit the axis regularity of the data, provided conditions (AR1)–(AR3) hold.
@@
Step 1: Corner geometry. Near a corner point $\bar{p}\in\bar{\Sigma}\cap\partial\mathcal{Q}$, introduce local coordinates $(\sigma,\tau)$ where $\sigma\geq 0$ is the distance to the axis $\partial\mathcal{Q}$ and $\tau$ is arc-length along $\partial\mathcal{Q}$, with $\tau=0$ at the corner. The MOTS curve $\bar{\Sigma}$ meets $\partial\mathcal{Q}$ at an angle $\theta_{0}\in(0,\pi)$, so near the corner:
@@ 4.3 Axisymmetric Setting
The key perturbation estimate underlying Theorem 4.13 is $|\mathcal{T}[f]|=O(s)$ as $s\to 0$. Since the principal Jang operator terms scale as $O(s^{-1})$ near the MOTS, the twist is subdominant by a factor of $s^{2}$. This decay follows from three facts: the twist 1-form satisfies $|\omega|\leq C_{\omega,\infty}$ by elliptic regularity, the $\rho^{2}$ prefactor remains bounded, and the gradient normalization $\sqrt{1+|\nabla f|^{2}}=O(s^{-1})$ from the Jang blow-up. Crucially, this estimate holds in the weighted Hölder spaces $C^{k,\beta}_{\rho}$ adapted to the axis geometry (see Appendix H), ensuring that the perturbation closes without regularity loss at the poles.
@@
$\mathcal{J}_{\text{axi}}[\bar{f}]$ $:=\bar{g}^{ij}\left(\frac{\bar{\nabla}_{ij}\bar{f}}{\sqrt{1+|\bar{\nabla}\bar{f}|^{2}}}-\bar{K}_{ij}\right)$
$\quad-\frac{\bar{f}^{i}\bar{f}^{j}}{1+|\bar{\nabla}\bar{f}|^{2}}\left(\frac{\bar{\nabla}_{ij}\bar{f}}{\sqrt{1+|\bar{\nabla}\bar{f}|^{2}}}-\bar{K}_{ij}\right)-\mathcal{T}[\bar{f}],$
@@
Let $(M^{3},g,K)$ be asymptotically flat, axisymmetric initial data with a stable outermost MOTS $\Sigma$. Let $s=\mathrm{dist}(\cdot,\Sigma)$ denote the signed distance to $\Sigma$, and let $\mathcal{T}[f]$ be the twist perturbation term (11). Then there exist constants $C_{\mathcal{T}}>0$ and $s_{0}>0$ depending only on the initial data such that
@@
Explicitly, $C_{\mathcal{T}}=C_{\omega,\infty}\cdot\rho_{\max}^{2}/C_{0}$, where $C_{\omega,\infty}=\sup_{\mathcal{Q}}|\omega|$, $\rho_{\max}=\sup_{\Sigma}\rho$, and $C_{0}=|\theta^{-}|/2$ is the Jang blow-up coefficient. The constant $C_{\mathcal{T}}$ depends only on the initial data and MOTS geometry, not on the Jang solution being constructed, so the perturbation argument is not circular.